# Physics of droplet regulation in biological cells

### v1

@@ 5.3.4  Spontaneous division
droplets grow by taking up material from the surrounding supersaturated dilute phase. If droplets are slightly deformed from a spherical state, e.g., due to thermal fluctuations, bumps on their surface have larger curvature and thus receive a higher material influx. This amplifies these bumps, leading to an instability similar to the Mullins-Sekerka instability of growing fronts [428]. Externally maintained droplets additionally lose material within their bulk, which reduces their total volume and causes a self-constriction in conjunction with the uneven influx. As a result, externally maintained droplets can divide spontaneously [139]. This shape instability can set in for droplets smaller than the stationary radius $R_{*}$ of isolated droplets predicted by equation (5.10b). Starting with an otherwise empty, and sufficiently large, system, a single droplet will thus grow and split multiple times until the entire system is filled with droplets. These duplications cease once droplets are so dense that they compete for material in the dilute phase. This competition lowers the effective supersaturation $\mathcal{S}$ and thus $R_{*}$, so the system finally settles in a stationary state with many droplets [139]. Spontaneous droplet division can also be explained using the electrostatic analogy since charged droplets also split in a similar fashion due to the Rayleigh instability [283, 429].
@@ 5.4 Internally maintained droplets
Internally maintained droplets of radius $R$ generally exhibit an efflux $J\approx 4\pi DR\left(\phi^{\mathrm{eq}}_{\mathrm{in}}-\phi_{\infty}\right)$, which follows from solving equation (5.8) in the stationary state; see figure 5.3B. Here, $\phi^{\mathrm{eq}}_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}(1+\ell_{\gamma}/R)$ is the equilibrium volume fraction inside the droplet and ${\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}=\phi^{(0)}_{\mathrm{out}}(1+\mathcal{S})}$ denotes the volume fraction in the dilute phase. To nucleate a droplet, the efflux $J$ needs to be compensated by local production of droplet material, which could originate from localised enzymes, e.g., at the surface of a catalytically active core of radius $a$. Since the integrated reaction flux $Q$ needs to exceed $J$, there is a minimal flux $Q_{\mathrm{min}}$ beyond which an internally maintained droplet is nucleated. For weak surface tension effects ($\ell_{\gamma}\ll a$), we find $Q_{\mathrm{min}}\approx 4\pi Da(\phi^{(0)}_{\mathrm{in}}-\phi_{\infty})$, whereas strong surface tension ($\ell_{\gamma}\gg a$) implies $Q_{\mathrm{min}}\approx 4\pi D\ell_{\gamma}\phi^{(0)}_{\mathrm{out}}$. In both cases, nucleation of internally maintained droplets is ensured, while spontaneous nucleation anywhere else in the system is virtually impossible.
If internally maintained droplets exceed $R_{\mathrm{crit}}$ given by equation (5.17), they grow even without catalytic cores, and their final size is only limited by the system size. Since larger droplets produce more droplet material, an emulsion of multiple internally maintained droplets is unstable, and coarsening is faster than the Ostwald ripening discussed in section 2.2.4 [383, 400]. Interestingly, however, sufficiently strong catalytically active cores can suppress Ostwald ripening, so the final droplet volumes approach values proportional to their respective reaction fluxes $Q$ [383]. The diffusive fluxes created catalytically active cores also automatically centre these cores within the droplets [403]. Taken together, these properties allow exquisite control over the positioning of internally maintained droplets [220].
@@ 5.5 More complex chemically active droplets
The behaviour of chemically active phase-separating systems can be much more complex when more species are involved. This topic represents a relatively new field of research, and so we only list a few recent discoveries here.
@@ 5.5.2  Self-organized drift
The combination of chemical turnover and diffusion generically creates concentration gradients, which can move particles and droplets by diffusiophoresis [431, 402, 178, 179, 432]. Gradients also develop around chemically active droplets (see figure 5.3), implying that they influence other droplets. For instance, externally maintained droplets naturally repel each other, which stabilises their hexagonal stationary pattern. Their repulsion follows directly from the electrostatic analogy discussed in section 5.3.2.
Chemical gradients also induce surface tension gradients, and thus movement via Marangoni flows, as discussed in section 3.3. While droplets generically react to external gradients, there is also the exciting possibility that they may create their own gradients. For instance, they can exploit an instability where movement amplifies the gradient, so the entire system settles in a moving state, either via diffusiophoresis [402] or Marangoni flows [267, 64]. Alternatively, chemical reactions can create an asymmetric state, which then moves due to isotropies in surrounding gradients [178]. In all these cases, activity needs to exceed a threshold, above which isotropy is spontaneously broken and swimming begins.
@@ 5.5.3  Active multi-component mixtures
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [434, 393, 394, 420] to systems with multiple length scales [435], to more complex spatiotemporal behaviour [436, 437]. Particularly interesting is the case where phase separation interferes with chemical reactions [438, 439], so that behaviours beyond those described above can emerge. For instance, droplets can now be stable even in the presence of passive reactions, which is impossible in binary systems [390]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [399, 383]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [406]. These droplets can also assume shell-like structures [401, 440], essentially because chemical reactions deplete the inside of the droplet so that concentration (shown in figure 5.3a) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [441] or they are fixed in particular locations [442]. As this field becomes more developed, there will surely be many additional phenomena that we are yet unveil.
@@ 6.1 Life cycle of droplets in biological cells
Droplets generally grow by taking up material from their surrounding if they are not otherwise restricted. Growth thus naturally stops once the surrounding environment is no longer supersaturated. However, growth may be limited if droplets are charged (section 3.1), the surrounding provides an elastic resistance (section 4.4.1), or chemical reactions resist further growth (section 5.3.1). Interestingly, all three of these mechanisms can be conceptualised as long-ranged repulsion due to electrostatics, elastic deformation, and reaction-diffusion systems, respectively. These long-ranged repulsive forces oppose the short-ranged attraction due to phase separation, resulting in a preferentially selected droplet size [416].
@@
Droplets need to be dissolved when they are no longer needed [28]. Cells can achieve dissolution by breaking down the involved biomolecules into their basic constituents to lower concentrations below saturation, so droplets spontaneously dissolve. This process is costly, time-consuming, and difficult to revert, but it removes droplets permanently. Alternatively, active reactions can be used to modify physical properties of the involved molecules, thus achieving faster, cheaper, and more flexible dissolution, which might only be temporary.
@@ 6.2 Droplet–environment interactions
Droplets react sensitively to their environment. Changes in external parameters, such as temperature, pH, and cellular composition, affect whether droplets form and how large they can grow (section 2.3.1). This ability to sense environmental cues allows droplets to react to changes in the internal states of cells, such as stress [451], which can then trigger appropriate responses. In addition, the first-order phase transition associated with phase separation implies hysteresis; droplets will not dissolve easily once formed, but also will not re-form easily once dissolved. This switch-like behaviour limits how much disruption can be caused by noise, allowing cells to make committed decisions [452, 453, 454].
@@
The interplay of droplets with each other and their environment leads to complex interaction networks, which can process information in cells. Such interaction networks typically involve multiple feedback loops, which enable additional dynamic behaviours [456]. At an abstract level, droplets might thus play an integral part in information processing and learning in biological cells [311].
@@ A Compressible multi-component fluid
$w_{i}$ $=\frac{1}{k_{\mathrm{B}}T}\left[\frac{e_{i}}{\Omega_{i}}-\frac{e_{0}}{\Omega_{0}}+\frac{e_{i0}+e_{0i}}{\nu_{0}\Omega_{i}}-\frac{2e_{00}}{\nu_{0}}\right]$ (A.9a)
@@ C Interface properties of binary liquids
$\gamma$ $=\kappa\int_{\phi^{(0)}_{\mathrm{out}}}^{\phi^{(0)}_{\mathrm{in}}}\phi_{\mathrm{eq}}^{\prime}\text{d}\phi_{\mathrm{eq}}$ (C.8)
@@ D Droplets in two dimensions
Droplet growth is still described by equation (2.36), but the diffusive fluxes $\boldsymbol{j}$ are more difficult to determine in two dimensions since steady state concentration profiles diverge logarithmically. As a result, one must introduce a cut-off length in numerical simulations [140] or use asymptotic matching [466, 467]. The latter approach shows that Ostwald ripening is suppressed [466] and that droplets drift in gradients [467], similar to their 3D counterparts. The diffusivity $D_{\mathrm{drop}}$ associated with Brownian motion also exhibits a logarithmic scaling. For a droplet of radius $R$ embedded in a membrane of thickness $h$, the Saffman-Delbrück model predicts [468]
@@
where $\gamma_{\mathrm{EM}}\approx 0.5772$ is the Euler-Mascheroni constant and $\eta_{\mathrm{m}}$ and $\eta_{\mathrm{b}}$ are the viscosities of the membrane and the surrounding bulk liquid, respectively.
@@ E Derivation of equilibrium stress tensor
from the surface contributions. To obey translational invariance, the free energy must be constant ($\delta F=0$) for a uniform displacement ($u_{\beta}=\text{const}$). Using Gauss’s theorem, this yields the Gibbs-Duhem relation
@@ F Transition state theory
The rate of the chemical conversion between reactants $R$ and products $P$ depends on their thermodynamic state described by the respective chemical potentials $\mu_{R}$ and $\mu_{P}$, as well as kinetic details of the reaction $R\rightleftharpoons P$. Typically, reactions follow a reaction path that involves an energetic barrier limiting the rate. This barrier can be conceptualised by a transition state $T$, leading to the extended reaction [386, 387, 388, 389, 390]
@@
$s_{R\rightarrow P}$ $\approx\frac{s_{T}}{2}e^{\beta(\mu_{R}-\mu_{T})}$ (F.3a)
@@
Typically, the transition state physically resembles both the reactants and the product, so a simple model for its chemical potential $\mu_{T}$ is
@@
$s=k\left(e^{\beta\alpha\mu}-e^{\beta(\alpha-1)\mu}\right)\;,$ (F.5)
where $\mu=\mu_{R}-\mu_{P}$ is the thermodynamic driving force of the reaction, and $k=\frac{1}{2}s_{T}e^{w_{T}}$ can be interpreted as a kinetic pre-factor, which can depend on composition. The symmetric choice $\alpha=\frac{1}{2}$ then leads to equation (5.4). Note that the reaction fluxes used in refs. [390, 396, 87, 83] follow when the chemical potential of the transition state is constant ($\mu_{T}=w_{T}$).

### v2

@@ 5.3.4  Spontaneous division
droplets grow by taking up material from the surrounding supersaturated dilute phase. If droplets are slightly deformed from a spherical state, e.g., due to thermal fluctuations, bumps on their surface have larger curvature and thus receive a higher material influx. This influx then amplifies these bumps, leading to an instability similar to the Mullins-Sekerka instability of growing fronts [437]. Additionally, externally maintained droplets lose material within their bulk, which reduces their total volume and causes a self-constriction in conjunction with the uneven material influx. As a result, externally maintained droplets can divide spontaneously [139]. This shape instability can set in for droplets smaller than the stationary radius $R_{*}$ of isolated droplets predicted by equation (5.11b). Starting with an otherwise empty, and sufficiently large, system, a single droplet will thus grow and split multiple times until the entire system is filled with droplets. These duplications cease once droplets are so dense that they compete for material in the dilute phase. This competition lowers the effective supersaturation $\mathcal{S}$ and thus $R_{*}$, until the system finally settles into a stationary state with many droplets [139]. Spontaneous droplet division can also be explained using the electrostatic analogy since charged droplets split in a similar fashion due to the Rayleigh instability [288, 438].
@@ 5.4 Internally maintained droplets
Internally maintained droplets of radius $R$ exhibit an efflux of droplet material $J\approx 4\pi DR\left(\phi^{\mathrm{eq}}_{\mathrm{in}}-\phi_{\infty}\right)$ (see figure 5.3B), whose functional form follows from solving equation (5.9) in the stationary state. Here, $\phi^{\mathrm{eq}}_{\mathrm{in}}=\phi^{(0)}_{\mathrm{in}}(1+\ell_{\gamma}/R)$ is the equilibrium fraction inside the droplet and $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}=\phi^{(0)}_{\mathrm{out}}(1+\mathcal{S})$ denotes the volume fraction in the dilute phase. To nucleate a droplet, the efflux $J$ needs to be compensated for by local production of droplet material, which could originate from localised enzymes, e.g., at the surface of a catalytically active core of radius $a$. Since the integrated reaction flux $Q$ at to such a core needs to exceed $J$, there is a minimal flux $Q_{\mathrm{min}}$ beyond which an internally maintained droplet is nucleated. For weak surface tension effects ($\ell_{\gamma}\ll a$), we find $Q_{\mathrm{min}}\approx 4\pi Da(\phi^{(0)}_{\mathrm{in}}-\phi_{\infty})$, whereas strong surface tension ($\ell_{\gamma}\gg a$) implies $Q_{\mathrm{min}}\approx 4\pi D\ell_{\gamma}\phi^{(0)}_{\mathrm{out}}$. In both cases, nucleation of internally maintained droplets around the core is ensured, while spontaneous nucleation anywhere else in the system is virtually impossible.
If internally maintained droplets exceed the radius $R_{\mathrm{crit}}$ given by equation (5.18), they grow even in the absence of a catalytic core, and their final size is only limited by the system size. Since larger droplets produce more droplet material, an emulsion of multiple internally maintained droplets is unstable, and coarsening is faster than that exhibited in passive Ostwald ripening [408, 409]. Interestingly, however, sufficiently strong catalytically active cores can suppress Ostwald ripening, such that final droplet volumes approach values proportional to the respective reaction fluxes $Q$ [392]. The diffusive fluxes created by catalytically active cores also automatically centre these cores within droplets [412]. Taken together, these properties allow exquisite control over the positioning of internally maintained droplets [224].
@@ 5.5 More complex chemically active droplets
The behaviour of chemically active phase-separating systems can be much more complex than that described above when more species are involved. Since this topic represents a relatively new field of research, we list only a few recent discoveries here.
@@ 5.5.2  Self-organized drift
The combination of chemical turnover and diffusion generically creates concentration gradients, which can move particles and droplets by diffusiophoresis [440, 411, 179, 180, 441]. Concentration gradients also develop around chemically active droplets (see figure 5.3), implying that they can influence the positioning of other droplets. For instance, externally maintained droplets naturally repel each other, which stabilises their hexagonal stationary pattern. Their repulsion follows directly from the electrostatic analogy discussed in section 5.3.2.
Chemical gradients also induce surface tension gradients, which can drive movement via Marangoni flows, as discussed in section 3.3. While droplets generically react to external gradients, there is also the exciting possibility that they may create their own gradients. Moreover, they may exploit an instability where movement amplifies this gradient, so that the entire system settles into a moving state, either via diffusiophoresis [411] or Marangoni flows [272, 64]. Alternatively, chemical reactions can create an asymmetric state, which then moves due to isotropies in surrounding gradients [179]. In all these cases, activity needs to exceed a threshold, above which isotropy is spontaneously broken and swimming begins.
@@ 5.5.3  Active multi-component mixtures
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [443, 402, 403, 429] to systems with multiple length scales [444], to more complex spatiotemporal behaviour [445, 446]. Particularly interesting is the case where phase separation interferes with chemical reactions [447, 448], so that behaviours beyond those described above can emerge. For instance, droplets may now be stable even in the presence of passive reactions, which is impossible in binary systems [399]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [408, 392]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [415, 449]. In both of these cases, the droplet size distribution relaxes to a monodisperse (and universal) distribution at late times [449]. Externally maintained droplets can also assume shell-like structures [410, 450], essentially because chemical reactions deplete the inside of the droplet so that concentration (see figure 5.3a) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [451] or they are fixed in particular locations [452]. Recent work has also shown that enzymes themselves may form phase-separated clusters with a characteristic equilibrium size [453], similar to the size control exhibited by externally maintained droplets discussed in section 5.3.1. As this field becomes more developed, there will surely be many additional phenomena that we are yet unveil.
@@ 6.1 Life cycle of droplets in biological cells
Droplets generally grow by taking up material from their surrounding if they are not otherwise restricted. Growth thus naturally stops once the surrounding environment is no longer supersaturated. However, growth may be limited if droplets are charged (section 3.1), the surrounding provides an elastic resistance (section 4.4.1), or chemical reactions restrict further growth (section 5.3.1). Interestingly, all three of these mechanisms can be conceptualised as long-ranged repulsion due to electrostatics, elastic deformation, and reaction-diffusion systems, respectively. These long-ranged repulsive forces oppose the short-ranged attraction due to phase separation, resulting in a preferentially selected droplet size [425].
@@
Droplets need to be dissolved when they are no longer needed [28]. Cells can achieve dissolution by breaking down the involved biomolecules into their basic constituents to lower concentrations below saturation, so that droplets spontaneously dissolve. This process is costly, time-consuming, and difficult to revert, but it removes droplets permanently. Alternatively, active reactions can be used to modify physical properties of the involved molecules, thus achieving faster, cheaper, and more flexible dissolution, which might only be temporary.
@@ 6.2 Droplet–environment interactions
Droplets react sensitively to their environment. Changes in external parameters, such as temperature, pH, and cellular composition, affect whether droplets form and how large they can grow (section 2.3.1). This ability to sense environmental cues allows droplets to react to changes in the internal states of cells, such as stress [462], which can then trigger appropriate responses. In addition, the first-order phase transition associated with phase separation implies hysteresis: droplets will not dissolve easily once formed, but also will not re-form easily once dissolved. This switch-like behaviour limits how much disruption can be caused by noise, allowing cells to make committed decisions [463, 464, 465].
@@
The interplay of droplets with each other and their environment leads to complex interaction networks, which can process information in cells and enable communication between droplets [467]. Such interaction networks typically involve multiple feedback loops, which give rise to additional dynamic behaviours [468]. At an abstract level, droplets might thus play an integral part in information processing and learning in biological cells [318].
@@ A Compressible multi-component fluid
$w_{i}$ $=\frac{1}{k_{\mathrm{B}}T}\left[\frac{e_{i}}{\Omega_{i}}-\frac{e_{0}}{\Omega_{0}}+\frac{e_{i0}+e_{0i}}{\nu_{0}\Omega_{i}}-\frac{2e_{00}}{\nu_{0}}\right]\;,$ (A.9a)
@@ C Interface properties of binary liquids
$\gamma$ $=\kappa\int_{\phi^{(0)}_{\mathrm{out}}}^{\phi^{(0)}_{\mathrm{in}}}\phi_{\mathrm{eq}}^{\prime}\text{d}\phi_{\mathrm{eq}}\;,$ (C.8)
@@ D Droplets in two dimensions
Droplet growth is still described by equation (2.36), but the diffusive fluxes $\boldsymbol{j}$ are more difficult to determine in two dimensions since steady state concentration profiles diverge logarithmically. As a result, one must introduce a cut-off length in numerical simulations [140] or use asymptotic matching [478, 479]. The latter approach shows that Ostwald ripening is suppressed [478] and that droplets drift in gradients [479], similar to their 3D counterparts. The diffusivity $D_{\mathrm{drop}}$ associated with Brownian motion also exhibits a logarithmic scaling. For a droplet of radius $R$ embedded in a membrane of thickness $h$, the Saffman–Delbrück model predicts [480]
@@
where $\gamma_{\mathrm{EM}}\approx 0.5772$ is the Euler–Mascheroni constant and $\eta_{\mathrm{m}}$ and $\eta_{\mathrm{b}}$ are the viscosities of the membrane and the surrounding bulk liquid, respectively.
@@ E Derivation of equilibrium stress tensor
from the surface contributions. To obey translational invariance, the free energy must be constant ($\delta F=0$) for a uniform displacement ($u_{\beta}=\text{const}$). Using Gauss’s theorem, this yields the Gibbs–Duhem relation
@@ F Transition state theory
The rate of the chemical conversion between reactants $R$ and products $P$ depends on their thermodynamic state described by the respective chemical potentials $\mu_{R}$ and $\mu_{P}$, as well as the kinetic details of the reaction $R\rightleftharpoons P$. Typically, reactions follow a reaction path that involves an energetic barrier limiting the rate. This barrier can be conceptualised by a transition state $T$, leading to the extended reaction [395, 396, 397, 398, 399]
@@
$s_{R\rightarrow P}$ $\approx\frac{s_{T}}{2}e^{\beta(\mu_{R}-\mu_{T})}\;,$ (F.3a)
@@
The reaction fluxes used in refs. [399, 405, 87, 83] then follow when the chemical potential of the transition state is constant ($\mu_{T}=w_{T}$). However, typical transition states physically resemble both the reactants and the product, so a simple model for its chemical potential $\mu_{T}$ is
@@
$s=k\left(e^{-\beta\alpha\mu}-e^{\beta(1-\alpha)\mu}\right)\;,$ (F.5)
where $\mu=\mu_{P}-\mu_{R}$ is the thermodynamic driving force of the reaction, and $k=\frac{1}{2}s_{T}e^{\beta w_{T}}$ can be interpreted as a kinetic pre-factor, which can depend on composition.
A simple model for the rate $k$ follows from considering ideal solutions, where $\mu_{i}=\beta^{-1}(\ln\phi_{i}+w_{i})$. Demanding that $s$ reduces to mass action kinetics,
$s=k_{\rightarrow}\prod_{i=1}^{N_{\mathrm{c}}}\phi_{i}^{\sigma_{\rightarrow,i}}-k_{\leftarrow}\prod_{i=1}^{N_{\mathrm{c}}}\phi_{i}^{\sigma_{\leftarrow,i}}\;,$ (F.6)
with constant reaction rates $k_{\rightarrow}$ and $k_{\leftarrow}$, implies
[…]
where the first term ensures that the kinetics obey mass action kinetics in the ideal limit, whereas the second term captures the thermodynamic constraints of detailed balance of the rates.
The additional requirement of symmetry between reactants and products, i.e., that swapping reactants and products must turn $s$ into $-s$, implies $\alpha=\frac{1}{2}$, leading to equations (5.4) and (5.5).